Two point charges (+Q) and (-2Q) are fixed on the X-axis at position a and 2a from origin respectively.At what position on the axis ,the resultant electric field is zero?

Grade:12

1 Answers

Arun
6 years ago
 
–K2q/(a-x)+ Kq/x2 = 0 
 
2/ (a-x)2 = 1/ x2

x/(a-x) = 1/sqrt(2)
(a-x)/x = sqrt(2)
a/x - 1 = sqrt(2)
a/x = 1 + sqrt(2)
x = a/(1 + sqrt(2))
 
Frim origin,
Distance = 
a + x
 
Regards
Arun